OpenStudy (anonymous):

factor 7x^2n+40x^n-12 please :)

OpenStudy (tamtoan):

7 x (-12) = -84 = 42 x (-2) , and 42 + (-2) = 40 , can you figure out from here ?

OpenStudy (anonymous):

I think the equation not right!!!

OpenStudy (tamtoan):

@ayman171819, the polynomial is factorable

OpenStudy (anonymous):

I'm sorry I don't think I can figure it out :( I was absent the day we did this in class so I am more clueless than usual :(

OpenStudy (anonymous):

I'll try agin

OpenStudy (tamtoan):

let u = x^n ...so x^(2n) = (x^n)^2.. then you can rewrite that polynomial as 7u^2 + 40u - 12 = 7u^2 + 42u - 2u - 12 ...now can you take it from here?

OpenStudy (anonymous):

yes thankyou! :D

OpenStudy (anonymous):

wait.... actually I'm not sure what to do next? Why did you do 7u^2+42u-2u-12? Sorry

OpenStudy (tamtoan):

group the first 2 terms, then group the second 2 terms, see what is the common factor after that :)

OpenStudy (anonymous):

AH YOU'RE RIGHT! :D Thanks again :D

OpenStudy (tamtoan):

np, enjoy and good night :)